OpenStudy (anonymous):

The formula for a geometric sequence is an = 25(2) n − 1. What is the 3rd term of the sequence? help me please somebody

OpenStudy (anonymous):

Perhaps question should be [a_{n} = 25(2n-1) Given \[a_{n} = 25(2n-1)\] In order to find 3rd term of the given sequence put n=3 so we find \[a_{3} = 25(2n-1)=25(2\times3-1)=25(6-1)=25\times5=125\] therefore third term = 125

OpenStudy (anonymous):

Perhaps question should be \[a_{n} = 25(2n-1)\]

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

so is it 75

OpenStudy (anonymous):

The formula for a geometric sequence is an = 25(2) n − 1. What is the 3rd term of the sequence? Answer 50 150 75 100

OpenStudy (anonymous):

this is mii question

OpenStudy (amistre64):

let n=3 ... and work the math

OpenStudy (amistre64):

and it helps if you notate it correctly an = 25 2^(n-1)

OpenStudy (anonymous):

dnt gt it

OpenStudy (amistre64):

then you really need to go review the basics of your material .....